y=-3x write an equation of a line that is perpendicular to y=-3x+5 and passes through 4,3
Answer by stanbon(75887)

write an equation of a line that is perpendicular
to y=-3x+5 and passes through 4,3
----
The given line has slope = -3
-----
Any perpendicular line must have slope = 1/3
----------
Form: y = mx + b
Solve for "b":
3 = (1/3)4 + b
-----
b = 5/3
--------
Equation:
y = (1/3)x + (5/3)
---------------------
Cheers,
Stan H.
=====================